OpenStudy (anonymous):

differentiate ln(secx).... is it 1/secx • √tanx ?

OpenStudy (anonymous):

shoot i closed the question accidentally agh

OpenStudy (anonymous):

@hartnn heya :3 any idea?^^

OpenStudy (amistre64):

take the derivative? yeah

OpenStudy (anonymous):

was my answer correct?

OpenStudy (amistre64):

what is the derivative of ln(u) with respect to x ?

OpenStudy (anonymous):

1/secx • √tanx ?

OpenStudy (anonymous):

or is it just √tanx ?

OpenStudy (amistre64):

\[ln(u)\to \frac{u'}{u}\] u = tan(x)

OpenStudy (anonymous):

is there any way to do it without subbing an unknown?

OpenStudy (amistre64):

its not that its an unknown, but this is just an application of the chain rule

OpenStudy (anonymous):

but the chain rule doesn't need substitution...

OpenStudy (amistre64):

f(g) = f'(g) * g' f = ln(x), g=tan(x)

OpenStudy (amistre64):

the chain rule is present in all the rules youve covered. you just havent realized it

OpenStudy (anonymous):

i use the chain rule, its like one of the three rules, but i don't substitute anything ><

OpenStudy (amistre64):

the derivative of ln(x) is x'/x what is x' with repect to x tho? dx/dx = 1

OpenStudy (amistre64):

you are using a mental shortcut when you 'dont substitute' anything.

OpenStudy (anonymous):

isn't it 1/x?

OpenStudy (anonymous):

i mean, the formula they give me is basically dy/dx = d/dx (lnx) = 1/x

OpenStudy (amistre64):

only with resect to x what is x' with respect to time? dx/dt which isnt always going to equal 1

OpenStudy (anonymous):

why do you put a prime behind the x?

OpenStudy (amistre64):

its a simpler notation to me

OpenStudy (anonymous):

what does it mean?

OpenStudy (amistre64):

x' is the derivative of x with respect to something. x' w.r.t x = dx/dx = 1 x' w.r.t. t = dx/dt x' w.r.t. g = dx/dg etc...

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

just as y' is the derivative of y w.r.t something y' w.r.t x = dy/dx y' w.r.t t = dy/dt y' w.r.t y = dy/dy = 1 etc ...

OpenStudy (amistre64):

this is how implicit derivatives are worked

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

so chain rule: f(g) derives to f'(g) * g' f = ln(x), g=tanx f' = 1/tanx, g' = sec^2(x)

OpenStudy (anonymous):

d/dx(tanx)=sec^2x

OpenStudy (amistre64):

lol, forgot what g was

OpenStudy (amistre64):

ln(secx) f = ln(x) and g = secx

OpenStudy (anonymous):

so what would integrate ln(secx) be?

OpenStudy (amistre64):

sec tan ------ = tan sec

OpenStudy (anonymous):

so basically, integrate ln(secx) = integrate ln(sec^2(x))

OpenStudy (amistre64):

not quite ln(a^b) = b ln(a) so the derivative of ln(sec^2) = 2 ln(sec)

OpenStudy (amistre64):

that wasnt worded right

OpenStudy (amistre64):

the derivative of ln(sec^2) = the derivaitve of2 ln(sec)

OpenStudy (anonymous):

how do you integrate sec?

OpenStudy (amistre64):

sec was integrated long before integration was established :)

OpenStudy (anonymous):

whaaat but my question was how do you integrate ln(secx)?

OpenStudy (anonymous):

wait so its actually 1/2 tan?

OpenStudy (amistre64):

multiply by a useful form of 1 \[sec=\frac{sec^2+sectan}{sec+tan}\]

OpenStudy (amistre64):

im not sure we are on the same page here. what question are you asking at the moment?

OpenStudy (anonymous):

what is integrate ln(secx)?

OpenStudy (amistre64):

integrate ln(sec(x)) as opposed to the derivative of ln(sec(x)) right?

OpenStudy (anonymous):

sorry differentitate i mean ><

OpenStudy (amistre64):

whew, cuase the integration is ugly

OpenStudy (anonymous):

wow i'm confusing us all sorry xD

OpenStudy (amistre64):

lets take it in steps do you agree that we define the chain rule as: [f(g(x))] = f'(g(x)) times g'(x) ?

OpenStudy (anonymous):

okay hang on could you just write the answer here first :3

OpenStudy (amistre64):

i already did sec tan ------ = tan sec

OpenStudy (anonymous):

so differentiate ln(secx) is tanx, to confirm?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

therefore differentiate secx is secxtanx?

OpenStudy (amistre64):

g' = secx tanx correct

OpenStudy (anonymous):

what would differentiate sec^3x be then?

OpenStudy (amistre64):

(secx)^3 f = ()^3 and g = secx apply chain rule

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

can't i do sec^3x =3sec^2x(3secxtanx)?

OpenStudy (amistre64):

yes, but without the extra 3 in there

OpenStudy (anonymous):

oh?

OpenStudy (amistre64):

f' = 3()^2 g' = sec tan

OpenStudy (anonymous):

oh yes because when i take the derative i have to take simply secx right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

phew thank you :3

OpenStudy (amistre64):

youre welcome :)